OpenStudy (britneysmithleigh):

when x is divided by 5, the remainder is 4. when x is divided by 9, the remainder is 0. which of the following is a possible value for x

OpenStudy (britneysmithleigh):

I am dumb founded when it comes to the question. i know the answer is 144 i just need to know the steps that would help me to get 144.

geerky42 (geerky42):

144? Why?

geerky42 (geerky42):

What about 9? 9 / 5 = 1 r 4 9 / 9 = 1 r 0

geerky42 (geerky42):

You already have hint to begin with; x divided by 9, the remainder is 0. This means x is multiple of 9.

jhonyy9 (jhonyy9):

from what are given result x/5 =a+4 x/9 =b+0 << which of the following is a possible value for x >> where are ? why 144 is right ? 144/5 = 28 so 144 divided by 5 de remainder is 4 10 --- 44 40 ---- - 4 144/9 = 16 so 144 divided by 5 the remainder is zero 9 --- - 54 54 ---- 0 0 so from these above wrote result that 144 satisfie the properties of x what are given in the text of your exercise x=144 is the correct answer sure

geerky42 (geerky42):

9 is also correct. So why would we aim for bigger number?

OpenStudy (welshfella):

I guess 9 wasn't one of the choices.

OpenStudy (sshayer):

it is divisible by 9 so let number be of of the type 9t where t is an integer now 9t gives remainder 4 when divided by 5 so 9t-4 is divisible by 5 when t=1,number 9 when t=6,number=54 when t=11,number=99 when t=16,number=144 .............................................. add every time 5 to the value of t you get a number which satisfy both the conditions.

OpenStudy (mathstudent55):

The question includes "which of the following is a possible value for x" That makes me think you were given choices to pick from. 9 is an answer to the plain question, and so is 54, but if in the choices only 144 works, then that is the answer. When you post a question with choices, you need to post the choices too.